### v1

# Geodesics in the Deep Linear Network

Abstract.
We derive a general system of ODEs and associated explicit solutions in a special case for geodesics between full rank matrices in the deep linear network geometry. In the process, we characterize all horizontal straight lines in the invariant balanced manifold that remain geodesics under Riemannian submersion.
Key words and phrases:
Deep linear network, geodesics, Riemannian submersion
1. Introduction
Let $\mathbb{M}_{d}$, $\mathfrak{M}_{d}$, $\operatorname{Symm}_{d}$, $\mathbb{P}_{d}$, and $O_{d}$ denote the sets of all, invertible, symmetric, positive definite, and orthogonal $d\times d$ matrices respectively.
1.1. The Deep Linear Network.
The deep linear network (DLN) is a phenomenological model of the training dynamics of neural networks [1, 2, 6, 9]. Given a depth $N$ and width $d$, in its simplest form the DLN consists of the “upstairs” (high-dimensional) parameter space of NN $d\times d$ matrices $\mathbb{M}_{d}^{N}$ and a “downstairs” (low-dimensional) observable space $\mathbb{M}_{d}$, linked by the product map $\phi:\mathbb{M}_{d}^{N}\to\mathbb{M}_{d}$ that sends
(1.1) $\mathbf{W}=(W_{N},\ldots,W_{1})\mapsto X=W_{N}W_{N-1}\cdots W_{1}.$
Indeed, $\mathbb{M}_{d}^{N}$ has $Nd^{2}$ degrees of freedom while $\mathbb{M}_{d}$ only has $d^{2}$, modeling overparameterization without the complications of nonlinearities and bias terms in traditional neural networks. The DLN is used to explicitly analyze the interplay of overparameterization with gradient descent of an energy function $E:\mathbb{M}_{d}\to\mathbb{R}$. The gradient flow in parameter space
(1.2) $\dot{\mathbf{W}}=-\nabla_{\mathbf{W}}E(X)=-\nabla_{\mathbf{W}}E(\phi(\mathbf{W}))$
can be written out explicitly via the chain rule as
(1.3) $\dot{W}_{p}=-(W_{N}\dots W_{p+1})^{T}dE(X)(W_{p-1}\dots W_{1})^{T}$
for all $1\leq p\leq N$, where $dE(X)\in\mathbb{M}_{d}$ is given elementwise by $dE(X)_{ij}=\frac{\partial E}{\partial X_{ij}}$.
1.2. Balancedness and Riemannian Geometry
Interestingly, the DLN gradient flow (Equation 1.2) has invariant manifolds with associated metrics that illustrate the geometric structure underlying the DLN. One such invariant manifold is defined by the constraints
(1.4) $W_{p+1}^{T}W_{p+1}=W_{p}W_{p}^{T}$
for all $1\leq p\leq N-1$ [2, Theorem 1]. We further restrict to when all coordinates are rank $d$ and denote this manifold $\mathcal{M}$, the balanced manifold. When $\mathbf{W}\in\mathcal{M}$, $X=\phi(\mathbf{W})\in\mathfrak{M}_{d}$.
It is not hard to show that in $\mathcal{M}$, $W_{p}$ have the same singular values for all $1\leq p\leq N$ and the right singular vectors of $W_{p+1}$ align with the left singular vectors of $W_{p}$ for $1\leq p\leq N-1$. Thus, given the SVD of a matrix $X=Q_{N}\Sigma Q_{0}$, we can parameterize any $\mathbf{W}\in\mathcal{M}$ such that $\phi(\mathbf{W})=X$ using just $(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})$, where $\Lambda=\Sigma^{\frac{1}{N}}$ and $W_{p}=Q_{p}\Lambda Q_{p-1}^{T}$ are obtained by the SVD for all $1\leq p\leq N$. Therefore, there exists a bijective map
(1.5) $\begin{split}\xi:\mathbb{R}^{d}_{+}\times O_{d}^{N+1}&\to\mathbb{M}_{d}^{N},\\ (\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})&\mapsto(W_{N},W_{N-1},\ldots,W_{1}).\end{split}$
Because $\mathcal{M}$ is embedded within the flat $(\mathbb{M}_{d}^{N},F)$ where $F$ is the Euclidean metric, we can induce the natural metric $\iota$ on $\mathcal{M}$ from $F$. Downstairs, we can define a metric $g^{N}$ on $\mathfrak{M}_{d}$ as follows: for any $Z\in T_{X}\mathfrak{M}_{d}$,
(1.6) $\|Z\|_{g^{N}}^{2}=\operatorname{Tr}\left(Z^{T}\mathcal{A}_{N,X}^{-1}(Z)\right),$
where $\mathcal{A}_{N,X}:T_{X}\mathfrak{M}_{d}^{*}\to T_{X}\mathfrak{M}_{d}$ is the linear operator that maps from 1-forms to the tangent space with explicit form
(1.7) $Z\mapsto\sum_{p=1}^{N}\left(XX^{T}\right)^{\frac{N-p}{N}}Z\left(X^{T}X\right)^{\frac{p-1}{N}}.$
$g^{N}$ is crucial in the DLN because it can be used to express the training dynamics of the end-to-end matrix $X$ as a Riemannian gradient flow [3, Section 4]. Furthermore, these metrics in tandem provide a geometric perspective on the DLN model from [10, Theorem 8] and [11, Theorem 36], which we elaborate on in Section 3.4.1:
Theorem 1.1 ($\phi$ is a Riemannian Submersion).
The end-to-end product map $\phi:\mathcal{M}\to\mathfrak{M}_{d}$ is a Riemannian submersion from $(\mathcal{M},\iota)$ to $(\mathfrak{M}_{d},g^{N})$.
[5, Section 4] derive explicit formulas for geodesics in the Bures-Wasserstein geometry by identifying a Riemannian submersion structure from invertible matrices to the positive definite matrices. We present an alternative proof of the same geodesic formulas that also generalizes to the DLN, yielding the general ODEs for geodesic in $(\mathfrak{M}_{d},g^{N})$. Furthermore, Theorem 1.1 implies that the DLN also has a Riemannian submersion structure with $(\mathcal{M},\iota)$, $(\mathfrak{M}_{d},g^{N})$, and $\phi$, which we leverage to solve the general equations in a specific case.
2. Main Results
Our main results are a statement of ODEs for general geodesics in $(\mathfrak{M}_{d},g^{N})$ and explicit solutions when the singular vectors of the endpoints are related by a single matrix $Q\in O_{d}$. The solutions are obtained by characterizing horizontal straight lines in $\mathcal{M}$ in a manner similar to the classical theory of confocal quadrics [7].
Let $f_{\alpha}^{[1]}(\Lambda)$ where $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$ denote the matrix
(2.1) $\left(f_{\alpha}^{[1]}(\Lambda)\right)_{ij}=\begin{cases}\frac{\lambda_{i}^{\alpha}-\lambda_{j}^{\alpha}}{\lambda_{i}-\lambda_{j}}&i\neq j,\\ \alpha\lambda_{i}^{\alpha-1}&i=j.\end{cases}$
We discuss the origin of this matrix in Section $3.$
Theorem 2.1 (DLN Geodesic Equations).
$X(t)=U(t)\Sigma(t)V(t)^{T}$ is a geodesic between $A$ and $B$ on $(\mathfrak{M}_{d},g^{N})$ if it satisfies the Dirichlet boundary value problem
(2.2) $\begin{cases}\dot{X}=\mathcal{A}_{N,X}(P),\\ \dot{P}=-U\left(\sum_{p=1}^{N}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]\Sigma+\Sigma\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\right)V^{T},\\ X(0)=A,\\ X(1)=B,\end{cases}$
where $M_{1}$ and $M_{2}$ are the symmetric matrices
(2.3) $M_{1}$ $=(U^{T}PV)\Sigma^{\frac{2(p-1)}{N}}(U^{T}PV)^{T}$
(2.4) $M_{2}$ $=(U^{T}PV)^{T}\Sigma^{\frac{2(N-p)}{N}}(U^{T}PV).$
We discover these equations via direct computation in Hamiltonian coordinates (Section 3.3). Using Riemannian submersion, we are then able to derive explicit solutions to these equations for particular endpoints. As an important intermediate result, we characterize the straight lines in $\mathcal{M}$.
Lemma 2.2 (Characterization of Straight Lines on $\mathcal{M}$).
Let $\mathbf{A}=\xi(\Lambda,Q_{N},\ldots,Q_{0})$ and $\mathbf{B}=\xi(\tilde{\Lambda},\tilde{Q}_{N},\ldots,\tilde{Q}_{0})$ be the endpoints of interest. Let
(2.5) $R_{p}=\tilde{Q}_{p}^{T}Q_{p}\in O_{d}.$
Then, the line
(2.6) $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$
lies in the balanced manifold for all $t\in[0,1]$:
  1. (1)
    if $R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$.
  2. (2)
    if and only if
    (2.7) $(R_{p+1}-R_{p-1})M^{T}$
    are skew symmetric, where
    (2.8) $M=\tilde{\Lambda}^{-1}R_{p}\Lambda$
    for all $1\leq p\leq N-1$.
We are now ready to state the second main result. Let $A=U\Sigma V^{T}$ and $B=\tilde{U}\tilde{\Sigma}\tilde{V}^{T}$ be the endpoints in $\mathfrak{M}_{d}$ of the geodesic.
Theorem 2.3 (Explicit Formulas for DLN Geodesics).
If there exists $Q\in O_{d}$ such that $\tilde{U}=UQ$ and $\tilde{V}=VQ$,
  1. (1)
    there exists $\mathbf{A}\in\phi^{-1}(A)$ and $\mathbf{B}\in\phi^{-1}(B)$ such that $x(t)=(1-t)\mathbf{A}+t\mathbf{B}$ is a geodesic in $(\mathcal{M},\iota)$.
  2. (2)
    the geodesic in $(\mathfrak{M}_{d},g^{N})$ between $A$ and $B$ is given by
    (2.9) $\begin{split}X(t)=&\left((1-t)U\Sigma^{\frac{1}{N}}+t\tilde{U}\tilde{\Sigma}^{\frac{1}{N}}Q^{T}\right)\\ &\quad\left((1-t)\Sigma^{\frac{1}{N}}+tQ\tilde{\Sigma}^{\frac{1}{N}}Q^{T}\right)^{N-2}\\ &\quad\left((1-t)\Sigma^{\frac{1}{N}}V^{T}+tQ\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T}\right).\end{split}$
Remark 2.4.
Examples of when $A$ and $B$ satisfy this right-left rotation condition is if they are both symmetric or if they have the same singular vectors ($Q=\text{Id}$).
Remark 2.5.
When $A$ and $B$ are diagonal with positive entries (fractional powers are well defined), the geodesic simplifies to
(2.10) $\begin{split}X(t)&=\left((1-t)A^{1/N}+tB^{1/N}\right)^{N}\\ &=\sum_{p=0}^{N}\binom{N}{p}(1-t)^{p}t^{N-p}A^{p/N}B^{(N-p)/N}.\end{split}$
This simplification can be shown by applying the theorem and expanding the geodesic formulas using that diagonal matrices commute. The limiting behavior as $N\to\infty$ (recognizing $X(t)$ is the expectation of $A^{S}B^{1-S}$ where $S$ is the average of NN Bernoulli(1t)(1-t) random variables) is
(2.11) $X(t)=A^{1-t}B^{t}.$
We discuss this striking simplification more in Appendix A.
The two important tools we will use to prove Theorem 2.3 in Section 3.4 are (1) the characterization of straight lines on $\mathcal{M}$ (Lemma 2.2) and (2) an orthonormal basis for the horizontal space in $\mathcal{M}$ constructed by [10, 11, Theorem 34] (Lemma 3.10 and see [9] for an exposition). Straight lines are automatically geodesics on $(\mathcal{M},\iota)$ because $\iota$ is induced from the Euclidean metric on $\mathbb{M}_{d}^{N}$. We will then send a subset of these straight lines to new geodesics in $(\mathfrak{M}_{d},g^{N})$ to uncover the explicit formulas for geodesics as stated.
3. Proofs
3.1. Geodesics Review
Definition 3.1 (Geodesic).
A geodesic between $A,B\in\mathcal{M}$ on $(\mathcal{M},g)$ is an extrema of the action
(3.1) $\mathcal{A}[x]=\int_{0}^{1}g_{x(t)}(\dot{x}(t),\dot{x}(t))dt$
over the set of paths
(3.2) $\mathcal{S}=\{x:[0,1]\to\mathcal{M}|x(0)=A,x(1)=B,x\in\mathcal{C}^{1}([0,1])\}.$
The minimizing $x$ is known to satisfy the Euler-Lagrange equations
(3.3) $\frac{d}{dt}\frac{\partial L(x,\dot{x})}{\partial\dot{x}}=-\frac{\partial L(x,\dot{x})}{\partial x}.$
The dual of the Lagrangian optimization problem is the Hamiltonian formulation that reveals an equivalent set of equations of motion. Given a Lagrangian $L(x,\dot{x})$, the Hamiltonian is the Legendre transform of $L(x,\cdot)$
(3.4) $H(x,p)$ $=\sup_{\dot{x}}(\langle p,\dot{x}\rangle-L(x,\dot{x})).$
It is not hard to see that the change of variables is $p=\frac{\partial L}{\partial\dot{x}}$ when $\dot{x}$ is maximizing. The corresponding canonical equations to the Euler-Lagrange equations are
(3.5) $\dot{x}$ $=\partial_{p}H(x,p),$
(3.6) $\dot{p}$ $=-\partial_{x}H(x,p).$
3.2. The Bures-Wasserstein Geometry
The Bures-Wasserstein geometry [5] is a special case of the DLN model when $N=2$ and $W_{1}=W_{2}^{T}$. The parameter space becomes $\mathfrak{M}_{d}$ whereas the observable space is $\mathbb{P}_{d}$. The map $\pi:\mathfrak{M}_{d}\to\mathbb{P}_{d}$ is a Riemannian submersion that maps
$W\in\mathfrak{M}_{d}\mapsto X=WW^{T}\in\mathbb{P}_{d}.$
There are $d^{2}$ degrees of freedom in parameter space compared to only $\binom{d+1}{2}$ dimensions downstairs as $X$ is symmetric. The downstairs metric, now denoted $g^{\text{BW}}$, at a point $X\in\mathbb{P}_{d}$ reduces to a very simple form
(3.7) $g^{\text{BW}}(Z_{1},Z_{2})=\frac{1}{2}\operatorname{Tr}(Z_{1}^{T}\mathcal{L}^{-1}_{X}(Z_{2})),$
where $\mathcal{L}^{-1}_{X}(Z)$ is the Lyapunov operator that maps $Z$ to the matrix that solves
(3.8) $\mathcal{L}^{-1}_{X}(Z)X+X\mathcal{L}^{-1}_{X}(Z)=Z.$
It is also clear that $\mathcal{L}_{X}(Y)=XY+YX$. An important quantity in the Bures-Wasserstein geometry is the geometric mean of two matrices.
Definition 3.2 (Geometric Mean).
Given $A,B\in\mathbb{P}_{d}$, the geometric mean of $A$ and $B$ (denoted by A#BA\#B) is
(3.9) $A\#B=A^{1/2}(A^{-1/2}BA^{-1/2})^{1/2}A^{1/2}.$
Lemma 3.3 (Geometric Mean Properties).
We will use two key facts about the geometric mean, both of which can be proven directly from the definition.
  1. (1)
    A#BA\#B is the unique positive definite solution $M$ to the Riccati Equation
    (3.10) $B=MA^{-1}M.$
  2. (2)
    $A(A^{-1}\#B)=(AB)^{1/2}$ and $(A^{-1}\#B)A=(BA)^{1/2}$.
As aforementioned, the Bures-Wasserstein geometry acts as a motivating example for our results. Specifically, we present a derivation of the explicit formulas for geodesics by solving the Hamiltonian equations of motion directly that we will generalize to the DLN (although we will not solve the equations explicitly).
Theorem 3.4 (Bures-Wasserstein Geodesics).
The geodesic path $X:[0,1]\to\mathbb{P}_{d}$ between $A,B\in\mathbb{P}_{d}$ in $(\mathbb{P}_{d},g^{\text{BW}})$ satisfies the boundary value problem
(3.11) $\begin{cases}\dot{X}=\mathcal{L}_{X}(P),\\ \dot{P}=-P^{2},\\ X(0)=A,\\ X(1)=B,\end{cases}$
and has explicit form given by
(3.12) $X(t)=(1-t)^{2}A+t^{2}B+t(1-t)\left((AB)^{1/2}+(BA)^{1/2}\right).$
Proof of Theorem 3.4.
To begin, recall that $T_{X}\mathbb{P}_{d}\cong\text{Symm}_{d}$ and denote $Y=\dot{X}$. In order to construct the change of variable from Lagrangian into Hamiltonian coordinates, we must first compute $D_{Y}[L(X,Y)](H)$ where $H\in T_{X}\mathbb{P}_{d}$.
$D_{Y}\left[\frac{1}{2}\operatorname{Tr}(Y^{T}\mathcal{L}^{-1}_{X}(Y))\right](H)$ $=\frac{1}{2}\operatorname{Tr}(D_{Y}[Y^{T}\mathcal{L}^{-1}_{X}(Y)](H))$
$=\frac{1}{2}\operatorname{Tr}(H^{T}\mathcal{L}^{-1}_{X}(Y)+Y^{T}\mathcal{L}^{-1}_{X}(H))$
$=\frac{1}{2}(g_{X}(H,Y)+g_{X}(Y,H))$
(3.13) $=\operatorname{Tr}(\mathcal{L}^{-1}_{X}(Y)H)$
where the second to last equality holds because $H$ and $Y$ are in $T_{X}\mathbb{P}_{d}$ and the last equality follows from the cyclic property of trace. Thus, we obtain that $P=\mathcal{L}^{-1}_{X}(Y)$ and that $Y=\mathcal{L}_{X}(P)$. Now, we can compute the Hamiltonian using the cyclic property of trace and the definition of $P$ and $\mathcal{L}_{X}$ as
$H(X,P)$ $=\operatorname{Tr}(P^{T}\mathcal{L}_{X}(P))-\frac{1}{2}\operatorname{Tr}(\mathcal{L}_{X}(P)\mathcal{L}^{-1}_{X}(\mathcal{L}_{X}(P)))$
(3.14) $=\frac{1}{2}\operatorname{Tr}(P\mathcal{L}_{X}(P))=\operatorname{Tr}(P^{2}X).$
The system of differential equations is now clear from the form of $H(X,P)$. We can solve them to find exact formulas for $X$ and $P$. In particular, $\dot{P}=-P^{2}$ implies
$-P^{-1}\dot{P}P^{-1}$ $=I$
$\frac{d}{dt}(P^{-1})$ $=I$
(3.15) $P^{-1}(t)$ $=tI+P^{-1}_{0}.$
Thus, $P(t)=(tI+P^{-1}_{0})^{-1}$. Factoring out $P^{-1}_{0}$ from the interior term reveals that $P(t)$ is a time derivative:
$P(t)$ $=(P^{-1}_{0}(tP_{0}+I))^{-1}$
(3.16) $=(tP_{0}+I)^{-1}P_{0}$
(3.17) $=\frac{d}{dt}\log(tP_{0}+I).$
The first differential equation can be expanded as
(3.18) $\dot{X}=X(t)P(t)+P(t)X(t).$
Motivated by the similarity of this equation to the Lax equation, we make the ansatz for the solution form of $X$ as
(3.19) $X(t)=e^{C(t)}X_{0}e^{D(t)},$
which after plugging in and matching coefficients gives that $\dot{C}(t)=\dot{D}(t)=P(t)$ and implying that
(3.20) $C(t)=D(t)$ $=\int_{0}^{t}P(s)ds$
(3.21) $=\int_{0}^{t}\frac{d}{ds}\log(sP_{0}+I)ds$
(3.22) $=\log(tP_{0}+I).$
Substituting $C$ and $D$ back into the ansatz gives that
$X(t)$ $=\exp\{\log(tP_{0}+I)\}X_{0}\exp\{\log(tP_{0}+I)\}$
$=(tP_{0}+I)X_{0}(tP_{0}+I)$
(3.23) $=(tP_{0}+I)A(tP_{0}+I).$
Applying the final unused condition that $X_{1}=B$ helps find $P_{0}$:
(3.24) $B=(P_{0}+I)A(P_{0}+I),$
which from Lemma 3.3 implies that
(3.25) $P_{0}=(A^{-1}\#B)-I.$
Finally, substituting $P_{0}$ into our solution formula for $X(t)$, we conclude that
(3.26) $X(t)$ $=(t((A^{-1}\#B)-I)+I)A(t((A^{-1}\#B)-I)+I)$
(3.27) $=(t(A^{-1}\#B)+(1-t)I)A(t(A^{-1}\#B)+(1-t)I)$
(3.28) $=(1-t)^{2}A+t^{2}B+t(1-t)((A^{-1}\#B)A+A(A^{-1}\#B))$
(3.29) $=(1-t)^{2}A+t^{2}B+t(1-t)\left((AB)^{1/2}+(BA)^{1/2}\right).$
3.3. General Equations from Hamiltonian
We will now generalize the preceding calculation to the DLN. First, recall functional calculus on symmetric matrices [4, Chapter 2]. If $S=Q\Lambda Q^{T}$ is a symmetric matrix and $f:\mathbb{R}\to\mathbb{R}$ is a 1D functional, $f$ can be extended to the symmetric matrices as (abusing notation)
(3.30) $f(S)=Qf(\Lambda)Q^{T}=Q\text{diag}(f(\lambda_{1}),f(\lambda_{2}),\dots,f(\lambda_{d}))Q^{T}.$
Then, the differential of $f(S)$ at $S$ evaluated in the direction $Z$ is exactly
(3.31) $Df[S](Z)=Q\left(f^{[1]}(\Lambda)\circ\left[Q^{T}ZQ\right]\right)Q^{T}$
where $f^{[1]}$ is the special matrix given elementwise by the conditional
(3.32) $f^{[1]}(\Lambda)_{ij}=\begin{cases}\frac{f(\lambda_{i})-f(\lambda_{j})}{\lambda_{i}-\lambda_{j}}&i\neq j\\ \lim_{\lambda_{i}\to\lambda_{j}}\frac{f(\lambda_{i})-f(\lambda_{j})}{\lambda_{i}-\lambda_{j}}&i=j.\end{cases}$
$f^{[1]}_{\alpha}$ (Equation 2.1) now clearly comes from Equation 3.32 if $f(x)=x^{\alpha}$. We also have the following derivative computation.
Lemma 3.5 (Derivatives of $(XX^{T})^{\alpha}$).
Let $X=U\Sigma V^{T}$ be the singular value decomposition of $X$. Then,
(3.33) $D_{X}\left[\left(XX^{T}\right)^{\alpha}\right](Z)$ $=U\left[f_{\alpha}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]U^{T}.$
(3.34) $D_{X}\left[\left(X^{T}X\right)^{\alpha}\right](Z)$ $=V\left[f_{\alpha}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right]V^{T}.$
where $\tilde{Z}=U^{T}ZV$.
Proof.
We write out the proof for the first result - the second is similar. Notice that $XX^{T}$ is a symmetric matrix with diagonalization $U\Sigma^{2}U^{T}$. Thus, we apply functional calculus along with the chain rule to find that
(3.35) $D_{X}\left[\left(XX^{T}\right)^{\alpha}\right](Z)=U\left[f_{\alpha}^{[1]}\left(\Sigma^{2}\right)\circ\left(U^{T}\left(D_{X}\left[XX^{T}\right](Z)\right)U\right)\right]U^{T}.$
Clearly, we know that
(3.36) $D_{X}\left[XX^{T}\right](Z)=XZ^{T}+ZX^{T}.$
Rewriting $Z$ as $\tilde{Z}$ in the singular vector coordinates via the transformation $Z=U\tilde{Z}V^{T}$ nets the final simplification that gives the form in the Lemma statement. ∎
We also prepare another intermediate result that will be useful for algebraic manipulation.
Lemma 3.6.
Let $K$ and $M$ be symmetric matrices, $\Sigma$ be a diagonal matrix, and $A$ be an arbitrary matrix. Then, we have the following identities.
(3.37) $\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\operatorname{Tr}([K\circ(A\Sigma)]M)=\operatorname{Tr}(\Sigma[K\circ M]A),$
(3.38) $\operatorname{Tr}([K\circ(\Sigma A)]M)$ $=\operatorname{Tr}([K\circ(A^{T}\Sigma)]M)=\operatorname{Tr}([K\circ M]\Sigma A).$
Proof.
Again, we will just prove for the first equation and note the second follows from almost identical steps. Let’s first prove that the first expression is equal to the last. We will approach by expanding into coordinates, following Einstein notation convention where repeated indices are summed.
$\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=K_{ij}\sigma_{i}(A^{T})_{ij}M_{ji}$
$=\sigma_{i}K_{ij}M_{ji}A_{ji}$
$=\sigma_{i}K_{ij}M_{ij}A_{ji}$
(3.39) $=\operatorname{Tr}(\Sigma[K\circ M]A).$
Similarly, we can prove the equality of the second expression and the third expression using the symmetry of $K$.
$\operatorname{Tr}([K\circ(A\Sigma)]M)$ $=K_{ij}A_{ij}\sigma_{j}M_{ji}$
$=\sigma_{j}K_{ij}M_{ji}A_{ij}$
$=\sigma_{j}K_{ji}M_{ji}A_{ij}$
(3.40) $=\operatorname{Tr}(\Sigma[K\circ M]A).$
Following similar steps to the Bures Wasserstein case, we can find the Hamiltonian on the manifold $(\mathfrak{M}_{d},g^{N})$ to be
(3.41) $H(X,P)=\frac{1}{2}\operatorname{Tr}\left(P^{T}\mathcal{A}_{N,X}(P)\right).$
Using these coordinates, we will now obtain the general equations on the manifold $(\mathfrak{M}_{d},g^{N})$ through direct computation, proving Theorem 2.1.
Proof of Theorem 2.1.
This first differential equation follows from exactly the same computation as the Bures-Wasserstein case (Equation 3.2) but with $\mathcal{L}^{-1}$ replaced by $\mathcal{A}^{-1}_{N,X}$. The only nontrivial equation is the second equation and the associated computation of $D_{X}[H(X,P)](Z)$. We begin by applying Lemma 3.5.
(3.42) $D_{X}[H(X,P)](Z)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}\left(T_{U}+T_{V}\right),$
(3.43) $T_{U}=P^{T}U\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]U^{T}PV\Sigma^{\frac{2(p-1)}{N}}V^{T},$
(3.44) $T_{V}=P^{T}U\Sigma^{\frac{2(N-p)}{N}}U^{T}PV\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right]V^{T}.$
To clean up further, we also write $P$ in terms of the singular value coordinates by setting $\tilde{P}=U^{T}PV$, leaving just
(3.45) $T_{U}$ $=\tilde{P}^{T}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]\tilde{P}\Sigma^{\frac{2(p-1)}{N}}$
(3.46) $T_{V}$ $=\tilde{P}^{T}\Sigma^{\frac{2(N-p)}{N}}\tilde{P}\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right].$
Applying the trace identities from Lemma 3.6 to Equations 3.45 and 3.46, where we note $f^{[1]}_{\alpha}(\Sigma^{2})$ is symmetric, and substituting in $M_{1}$ and $M_{2}$ from Equation 2.3 reveal
[…]
(3.49) $=\operatorname{Tr}\left(V\sum_{p=1}^{N}\left(\Sigma\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]+\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\Sigma\right)U^{T}Z\right)$
This lets us conclude that the second canonical equation in the Hamiltonian formulation is
(3.50) $\dot{P}=-U\sum_{p=1}^{N}\left(\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]\Sigma+\Sigma\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\right)V^{T}.$
However, unlike the Bures-Wasserstein case (Theorem 3.4), we find that this final form is not particularly amenable to direct solution, despite the commutator structure and relation to the singular vector coordinates $U$ and $V$. Thus, we turn to special cases, where we will be able to find explicit formulas by leveraging the inherent geometric structure of the DLN model.
3.4. Explicit Solutions from Submersion
3.4.1. Riemannian Geometry Background
Recall the following important definitions and results from Riemannian geometry [5].
Definition 3.7 (Vertical and Horizontal Spaces).
Let $(\mathcal{M},g)$ and ($\mathcal{N},h)$ be linked by a differentiable map $\pi:\mathcal{M}\to\mathcal{N}$. Then, there exists a direct sum decomposition of the tangent space at any point $m\in\mathcal{M}$ via the differential of $\phi$.
(3.51) $T_{m}\mathcal{M}=\ker D\pi[m]\oplus(\ker D\pi[m])^{\perp}=\mathcal{V}_{m}\oplus\mathcal{H}_{m},$
where $\mathcal{V}_{m}$ and $\mathcal{H}_{m}$ denote the vertical and horizontal spaces at $m$ respectively.
Definition 3.8 (Riemannian Submersion).
Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be two Riemannian manifolds and let $T_{m}\mathcal{M}=\mathcal{V}_{m}\oplus\mathcal{H}_{m}$ be the direct sum decomposition of the tangent space at $m\in\mathcal{M}$. A mapping $\pi:\mathcal{M}\to\mathcal{N}$ is a Riemannian submersion if and only if the differential $D\pi[m]:T_{m}\mathcal{M}\to T_{\pi(m)}\mathcal{N}$ satisfies the following for all $m\in\mathcal{M}$:
  1. (1)
    $D\pi[m]$ is surjective ($\pi$ is a smooth submersion).
  2. (2)
    $D\pi[m]$ restricted to the horizontal space is an isometry i.e. $\|z\|_{g}=\|D\pi[m](z)\|_{h}$ for all $z\in\mathcal{H}_{m}$.
A core fact underlying [5, Section 4] and Theorem 2.3 is that understanding the horizontal space of a submersion allows us to identify geodesics that are preserved under the submersion.
Theorem 3.9 (Submersions Preserve Horizontal Geodesics).
Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be Riemannian manifolds and $\phi:\mathcal{M}\to\mathcal{N}$ be a Riemannian submersion. Let $x:[0,1]\to\mathcal{M}$ be a geodesic between $A,B\in\mathcal{M}$ in the geometry $(\mathcal{M},g)$ so that $x^{\prime}(0)\in\mathcal{H}_{A}$ i.e. is horizontal. Then,
  1. (1)
    $x^{\prime}(t)$ is horizontal for all $t\in[0,1]$.
  2. (2)
    $\phi\circ x$ is a geodesic in $(\mathcal{N},h)$ between $\phi(A)$ and $\phi(B)$ and is the same length as $x$.
Thus, if we are able to classify horizontal geodesics in $(\mathcal{M},\iota)$, Theorem 3.9 will automatically give explicit formulas for geodesics in $(\mathfrak{M}_{d},g^{N})$ through $\phi$, the submersion. The simplest class of geodesics in $(\mathcal{M},\iota)$ are straight lines. Thus, we begin with a proof of Lemma 2.2 that characterizes the straight lines in $\mathcal{M}$.
Proof of Lemma 2.2.
We wish to find the condition that $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$ is in $\mathcal{M}$ for all $t\in(0,1)$ if $\mathbf{A},\mathbf{B}\in\mathcal{M}$. Fix $p\in\{1,2,\ldots,N-1\}$. We would like to show that
(3.52) $W_{p+1}(t)^{T}W_{p+1}(t)=W_{p}(t)W_{p}(t)^{T},$
where $W_{p}(t)=(1-t)W_{p}(0)+tW_{p}(1)$ for all $p$. Expanding the algebra out and using that $W_{p}(0)=A_{p}$ and $W_{p}(1)=B_{p}$ are both in the balanced manifold, we find that $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in(0,1)$ if and only if
(3.53) $B_{p+1}^{T}A_{p+1}+A_{p+1}^{T}B_{p+1}=B_{p}A_{p}^{T}+A_{p}B_{p}^{T}.$
Now, we use the parameterizations of $A$ and $B$ to expand into the orthogonal matrix components.
(3.54) $\begin{split}\tilde{Q}_{p}\tilde{\Lambda}(\tilde{Q}_{p+1}^{T}Q_{p+1})\Lambda Q_{p}^{T}&+Q_{p}\Lambda(Q_{p+1}^{T}\tilde{Q}_{p+1})\tilde{\Lambda}\tilde{Q}_{p}^{T}\\ &=\tilde{Q}_{p}\tilde{\Lambda}(\tilde{Q}_{p-1}^{T}Q_{p-1})\Lambda Q_{p}^{T}+Q_{p}\Lambda(Q_{p-1}^{T}\tilde{Q}_{p-1})\tilde{\Lambda}\tilde{Q}_{p}^{T}.\end{split}$
Rewriting using $R_{p}$ notation and left and right multiplying by $\tilde{Q}_{p}^{T}$ and $Q_{p}$ respectively gives
(3.55) $\tilde{\Lambda}(R_{p+1}-R_{p-1})\Lambda$ $=-R_{p}\Lambda(R_{p+1}-R_{p-1})^{T}\tilde{\Lambda}R_{p}.$
The sufficient condition ($R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$) clearly results in this condition being satisfied for all $t\in(0,1)$. From here, it is a simple rewriting in terms of $M$ to get the final equivalent condition:
(3.56) $R_{p+1}-R_{p-1}$ $=-M(Q_{p+1}-Q_{p-1})^{T}M^{-T}.$
(3.57) $(R_{p+1}-R_{p-1})M^{T}$ $=-M(Q_{p+1}-Q_{p-1})^{T}.$
The structure demonstrates that the difference
(3.58) $R_{p+1}-R_{p-1}$
behaves like a skew symmetric matrix after changing into coordinates given by $M$ (or essentially $R_{p}$). ∎
We now state an orthonormal basis for the horizontal space in $\mathcal{M}$ computed in [10, Section 6] and [11, Theorem 34]. Let
(3.59) $\mathbf{M}\in\mathcal{M},\mathbf{M}=\left(Q_{N}\Lambda^{\frac{1}{N}}Q_{N-1}^{T},Q_{N-1}\Lambda^{\frac{1}{N}}Q_{N-2}^{T},\dots,Q_{1}\Lambda^{\frac{1}{N}}Q_{0}^{T}\right)$
so that $\mathbf{M}=\xi(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{1})$ and $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$. Furthermore, let $q_{p,i}$ denote the $i$th column of $Q_{p}$. Then, define the following special rank one elements of $\mathbb{M}_{d}^{N}$:
(3.60) $\begin{cases}\mathbf{l}^{k}=\left(l^{k}_{N},l^{k}_{N-1},\ldots,l^{k}_{1}\right),\\ l^{k}_{s}=\frac{1}{\sqrt{N}}q_{s,k}q_{s-1,k}^{T},\end{cases}$
(3.61) $\begin{cases}\mathbf{u}^{k,l,0}=\left(u_{N}^{k,l,0},u_{N-1}^{k,l,0},\ldots,u_{1}^{k,l,0}\right),\\ u^{k,l,0}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{s-1}\lambda_{l}^{N-s}q_{s,l}q_{s-1,k}^{T},\end{cases}$
and
(3.62) $\begin{cases}\mathbf{u}^{k,l,N}=\left(u_{N}^{k,l,N},u_{N-1}^{k,l,N},\ldots,u_{1}^{k,l,N}\right),\\ u^{k,l,N}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{N-s}\lambda_{l}^{s-1}q_{s,k}q_{s-1,l}^{T}.\end{cases}$
Lemma 3.10 (Orthonormal Basis for Horizontal Space in $\mathcal{M}$).
(3.63) $\begin{cases}\mathbf{l}=\{l_{k}|1\leq k\leq d\},\\ \mathbf{u}^{0}=\{u^{k,l,0}|1\leq k<l\leq d\},\\ \mathbf{u}^{N}=\{u^{k,l,N}|1\leq k<l\leq d\}\end{cases}$
form a basis for $\mathcal{H}_{\mathbf{M}}=(\ker\phi^{*}[\mathbf{M}])^{\perp}$.
Remark 3.11.
Note a matrix $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{M}}$ (span of $\mathbf{l}$, $\mathbf{u}^{k,l,0}$, and $\mathbf{u}^{k,l,N}$) if and only if there exists a set of coefficients $\alpha\in\mathbb{M}_{d}$ such that
(3.64) $Q_{p}\alpha Q_{p-1}^{T}=Z_{p}$
for all $1\leq p\leq N$.
Now that we understand both the straight lines and horizontal space in $\mathcal{M}$, we can use $\phi$ and Theorem 3.9 to construct new explicit formulas for geodesics on $\mathfrak{M}_{d}$.
Proof of Theorem 2.3.
It is sufficient to first construct $\mathbf{A}\in\phi^{-1}(A)\subseteq\mathcal{M}$ and $\mathbf{B}\in\phi^{-1}(B)\subseteq\mathcal{M}$ such that
(3.65) $x(t)=(1-t)\mathbf{A}+t\mathbf{B}\in\mathcal{M}\text{ for all }t\in[0,1]$
because $\iota$ is induced by the flat Frobenius metric. Then, we will show that $x^{\prime}(t)$ is horizontal at $t=0$ by proving that
(3.66) $\frac{d}{dt}\left((1-t)\mathbf{A}+t\mathbf{B}\right)=-\mathbf{A}+\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$
or that equivalently $\mathbf{A}\in\mathcal{H}_{\mathbf{A}}$ and $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ because $\phi^{*}$ is linear. Finally, Theorem 3.9 can be applied, implying that $\phi\circ x$ is a geodesic in the downstairs space between $\phi(\mathbf{A})$ and $\phi(\mathbf{B})$ and giving the explicit formula as desired.
We begin by identifying the correct parameterizations for the upstairs endpoints. Consider the following parameterization for $\mathbf{A}$:
(3.67) $\mathbf{A}=\xi\left(\Sigma^{\frac{1}{N}},U,I,\ldots,I,V\right)$
where $\xi$ is the map given in Equation 1.5 since $\mathbf{A}\in\mathcal{M}$. Clearly, $\phi(\mathbf{A})=A$. Furthermore, Lemma 3.10 gives a set of orthonormal basis vectors for the horizontal space $\mathcal{H}_{\mathbf{A}}$:
(3.68) $l^{k}_{s}$ $=\frac{1}{\sqrt{N}}\begin{cases}u_{k}e_{k}^{T}&s=N\\ e_{k}v_{k}^{T}&s=1\\ e_{k}e_{k}^{T}&\text{o/w}\end{cases},$
(3.69) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{s-1}{N}}\sigma_{l}^{\frac{N-s}{N}}\begin{cases}u_{l}e_{k}^{T}&s=N\\ e_{l}v_{k}^{T}&s=1\\ e_{l}e_{k}^{T}&\text{o/w}\end{cases},$
(3.70) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{N-s}{N}}\sigma_{l}^{\frac{s-1}{N}}\begin{cases}u_{k}e_{l}^{T}&s=N\\ e_{k}v_{l}^{T}&s=1\\ e_{k}e_{l}^{T}&\text{o/w}\end{cases}.$
By Remark 3.11, $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{A}}$ if and only if there exists $\alpha\in\mathbb{M}_{d}$ such that
(3.71) $U\alpha$ $=Z_{N},$
(3.72) $\alpha$ $=Z_{p},2\leq p\leq N-2,$
(3.73) $\alpha V^{T}$ $=Z_{1}.$
It is immediately obvious that $\mathbf{A}$ is in its own horizontal space $\mathcal{H}_{\mathbf{A}}$ using the coefficient matrix $\alpha=\Sigma^{\frac{1}{N}}$.
Now, also consider
[…]
where $Q=U^{T}\tilde{U}=V^{T}\tilde{V}$ by assumption. Clearly, $\phi(\mathbf{B})=B$. Furthermore, $\mathbf{A}$ and $\mathbf{B}$ satisfy the sufficient condition of Lemma 2.2 since $Q^{T}=\tilde{U}^{T}U=\tilde{V}^{T}V$ and $R_{p}$ is constant for all $p$. Thus, the line $(1-t)\mathbf{A}+t\mathbf{B}$ is indeed the geodesic between $\mathbf{A}$ and $\mathbf{B}$ in $(\mathcal{M},\iota)$.
Now we will show $x^{\prime}(0)$ is horizontal for the chosen $\mathbf{A}$ and $\mathbf{B}$. We have already shown that $\mathbf{A}$ is in $\mathcal{H}_{\mathbf{A}}$. By Equation 3.66, it is now sufficient to show that $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ as well. We thus return to solving the system of equations in Remark 3.11 with $\mathbf{Z}=\mathbf{B}$:
(3.75) $U\alpha$ $=\tilde{U}\tilde{\Sigma}^{\frac{1}{N}}Q^{T},$
(3.76) $\alpha$ $=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T},$
(3.77) $\alpha V^{T}$ $=Q\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T},$
and has the obvious solution given by the middle equations of $\alpha=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T}$ which is consistent with the others by definition of $Q$. Thus, $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$.
Since we have shown that both $\mathbf{B}$ and $\mathbf{A}$ are in $\mathcal{H}_{\mathbf{A}}$ and that the geodesic linking them is a straight line, it follows from Equation 3.66 that $x^{\prime}(0)$ is horizontal. Applying Theorem 3.9, we conclude that $\phi\circ x$ is a geodesic in $(\mathfrak{M}_{d},g^{N})$ of the same length as $x$. The explicit formula comes directly from applying $\phi$ to $x$ between our parameterized $\mathbf{A}$ and $\mathbf{B}$. ∎
4. Open Problems
The main follow up work is (1) solving the general case geodesic equations for arbitrary endpoints and (2) generalizing geodesic results to low rank (rank <d<d) matrices, not just $\mathcal{M}$, as the Riemannian metric and submersion structure exists in all ranks [3, 10, 11]. Once geodesics and distances are understood in the geometries underlying the DLN, one can consider reconnecting them to the vector field given by the gradient flow to make conclusions about training dynamics, similar to [8, Section 3.3].
Acknowledgements
This work is based on a chapter of the author’s undergraduate thesis supervised by Govind Menon (Brown University), to whom the author is grateful towards for the initial problem statement and continual feedback on the project and paper. The author is also thankful for advice from Austin Stromme (ENSAE/CREST) and Kathryn Lindsey (Boston College) at various stages of this work. This work was partially supported by NSF grant 2407055.
Appendix A Informal Note on Equation 2.11
The geodesic in Equation 2.11 also appears in another geometry [4] when the endpoints are positive definite diagonal matrices. In particular, we endow $\mathbb{P}_{d}$ with the trace metric: at a matrix $P\in\mathbb{P}_{d}$ and tangent vector $Z\in T_{P}\mathbb{P}_{d}$, we let
(A.1) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(P^{-1}ZP^{-1}Z).$
Consider $P$ positive definite and diagonal, and let $Z$ commute with $P$ (e.g. also diagonal). By the properties of trace,
(A.2) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(ZP^{-2}Z).$
Between any matrices $P_{0},P_{1}\in\mathbb{P}_{d}$, the geodesic is exactly $P_{0}^{1-t}P_{1}^{t}$. Interestingly, $g^{N}$ as $N\to\infty$ coincides with the trace metric under these assumptions. Consider a variant of $\mathcal{A}_{N,\cdot}$ with an additional $\frac{1}{N}$ factor in front so that a $N\to\infty$ limit exists. Then, indeed, if $P$ is diagonal and commutes with $Z$,
(A.3) $\mathcal{A}_{N,P}(Z)=\frac{1}{N}\sum_{p=1}^{N}(P^{2})^{\frac{N-p}{N}}Z(P^{2})^{\frac{p-1}{N}}=(P^{2})^{\frac{N-1}{N}}Z.$
The limit as $N\to\infty$ is just $P^{2}Z$. Thus, $\mathcal{A}^{-1}_{\infty,P}=P^{-2}Z$ and we conclude that
(A.4) $\|Z\|_{g^{\infty}}=\operatorname{Tr}(Z\mathcal{A}^{-1}_{\infty,P}Z)=\operatorname{Tr}(ZP^{-2}Z)=\|Z\|_{\operatorname{Tr}}.$
More generally, if $X\in\mathfrak{M}_{d}$ is symmetric, for all $Z\in T_{X}\mathfrak{M}_{d}$ where $X$ and $Z$ commute, we have that this agreement between $g^{\infty}$ and the trace metric. Since along the geodesic $X(t)$ between $A$ and $B$ $\dot{X}\in T_{X}\mathfrak{M}_{d}$ is always diagonal, we have agreement between the trace metric and DLN metric for the entire geodesic. Contrast this with the smallest nontrivial depth, $N=2$, where if $X$ is symmetric we know already that the DLN metric coincides with the Bures Wasserstein metric for symmetric tangent vectors.

### v2

@@
 We derive a general system of ODEs and associated explicit solutions in a special case for geodesics between full rank matrices in the deep linear network geometry.
-In the process, we characterize all horizontal straight lines in the invariant balanced manifold that remain geodesics under Riemannian submersion.
+In the process, we find horizontal straight lines in the invariant balanced manifold that remain geodesics under Riemannian submersion.
 Key words and phrases:
 Deep linear network, geodesics, Riemannian submersion
+2020 Mathematics Subject Classification:
+Primary
 1. Introduction
@@
 The deep linear network (DLN) is a phenomenological model of the training dynamics of neural networks [1, 2, 6, 11].
-Given a depth $N$ and width $d$, in its simplest form the DLN consists of the “upstairs” (high-dimensional) parameter space of NN $d\times d$ matrices $\mathbb{M}_{d}^{N}$ and a “downstairs” (low-dimensional) observable space $\mathbb{M}_{d}$, linked by the product map $\phi:\mathbb{M}_{d}^{N}\to\mathbb{M}_{d}$ that sends
+Given a depth $N\in\mathbb{Z}^{+}$ and width $d\in\mathbb{Z}^{+}$, the DLN in its simplest form consists of the “upstairs” (high-dimensional) parameter space of NN $d\times d$ matrices $\mathbb{M}_{d}^{N}$ and a “downstairs” (low-dimensional) observable space $\mathbb{M}_{d}$, linked by the product map $\phi:\mathbb{M}_{d}^{N}\to\mathbb{M}_{d}$ that sends
 (1.1) $\mathbf{W}=(W_{N},\ldots,W_{1})\mapsto X=W_{N}W_{N-1}\cdots W_{1}.$
@@
 1.2. Balancedness and Riemannian Geometry
-Interestingly, the DLN gradient flow (Equation 1.2) has invariant manifolds with associated metrics that illustrate the geometric structure underlying the DLN.
-One such invariant manifold is defined by the constraints
+Interestingly, the DLN gradient flow (Equation 1.2) has invariant manifolds with associated metrics that illustrate the underlying geometric structure.
+One such invariant manifold is defined by the balancedness constraints
 (1.4) $W_{p+1}^{T}W_{p+1}=W_{p}W_{p}^{T}$
 for all $1\leq p\leq N-1$ [2, Theorem 1].
-We further restrict to when all coordinates are rank $d$ and denote this manifold $\mathcal{M}$, the balanced manifold.
-When $\mathbf{W}\in\mathcal{M}$, $X=\phi(\mathbf{W})\in\mathfrak{M}_{d}$.
-It is not hard to show that in $\mathcal{M}$, $W_{p}$ have the same singular values for all $1\leq p\leq N$ and the right singular vectors of $W_{p+1}$ align with the left singular vectors of $W_{p}$ for $1\leq p\leq N-1$.
-Thus, given the SVD of a matrix $X=Q_{N}\Sigma Q_{0}$, we can parameterize any $\mathbf{W}\in\mathcal{M}$ such that $\phi(\mathbf{W})=X$ using just $(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})$, where $\Lambda=\Sigma^{\frac{1}{N}}$ and $W_{p}=Q_{p}\Lambda Q_{p-1}^{T}$ are obtained by the SVD for all $1\leq p\leq N$.
-Therefore, there exists a bijective map
+When all coordinates are also rank $d$, the manifold is called the balanced manifold and is denoted as $\mathcal{M}$.
+Note that when $\mathbf{W}\in\mathcal{M}$, $X=\phi(\mathbf{W})\in\mathfrak{M}_{d}$.
+In $\mathcal{M}$, $W_{p}$ have the same singular values for all $1\leq p\leq N$ and the right singular vectors of $W_{p+1}$ align with the left singular vectors of $W_{p}$ for $1\leq p\leq N-1$.
+Thus, given the SVD of a matrix $X=Q_{N}\Sigma Q_{0}^{T}$, we can parameterize any $\mathbf{W}\in\mathcal{M}$ such that $\phi(\mathbf{W})=X$ using just
+(1.5) $(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})\in\mathbb{R}_{+}^{d}\times O_{d}^{N+1},$
+where $\Lambda=\Sigma^{\frac{1}{N}}$ and $W_{p}=Q_{p}\Lambda Q_{p-1}^{T}$ are obtained by the SVD for all $1\leq p\leq N$. Therefore, there exists a bijective map
 (1.6) $\begin{split}\xi:\mathbb{R}^{d}_{+}\times O_{d}^{N+1}&\to\mathbb{M}_{d}^{N},\\ (\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})&\mapsto(W_{N},W_{N-1},\ldots,W_{1}).\end{split}$
@@
 $g^{N}$ is crucial in the DLN because it can be used to express the training dynamics of the end-to-end matrix $X$ as a Riemannian gradient flow [3, Section 4].
-Furthermore, these metrics in tandem provide a geometric perspective on the DLN model from [10, Theorem 8] and [11, Theorem 36], which we elaborate on in Section 3.4.1:
+Furthermore, these metrics in tandem provide a geometric perspective on the DLN model from [10, Theorem 8] and [13, Theorem 36], which we elaborate on in Section 3.4.1:
 Theorem 1.1 ($\phi$ is a Riemannian Submersion).
 The end-to-end product map $\phi:\mathcal{M}\to\mathfrak{M}_{d}$ is a Riemannian submersion from $(\mathcal{M},\iota)$ to $(\mathfrak{M}_{d},g^{N})$.
-[5, Section 4] derive explicit formulas for geodesics in the Bures-Wasserstein geometry by identifying a Riemannian submersion structure from invertible matrices to the positive definite matrices.
-We present an alternative proof of the same geodesic formulas that also generalizes to the DLN, yielding the general ODEs for geodesic in $(\mathfrak{M}_{d},g^{N})$.
+Explicit formulas for geodesics in the Bures-Wasserstein geometry can be derived by identifying a Riemannian submersion structure from invertible matrices to the positive definite matrices [4, Section 4].
+We present an alternate derivation of the geodesic formula that generalizes to the DLN, yielding the general ODEs for geodesics in $(\mathfrak{M}_{d},g^{N})$.
 Furthermore, Theorem 1.1 implies that the DLN also has a Riemannian submersion structure with $(\mathcal{M},\iota)$, $(\mathfrak{M}_{d},g^{N})$, and $\phi$, which we leverage to solve the general equations in a specific case.
@@
 Our main results are a statement of ODEs for general geodesics in $(\mathfrak{M}_{d},g^{N})$ and explicit solutions when the singular vectors of the endpoints are related by a single matrix $Q\in O_{d}$.
-The solutions are obtained by characterizing horizontal straight lines in $\mathcal{M}$ in a manner similar to the classical theory of confocal quadrics [7].
+The solutions are obtained by finding a sufficient condition for horizontal straight lines in $\mathcal{M}$ in a manner similar to the classical theory of confocal quadrics [8].
 Let $f_{\alpha}^{[1]}(\Lambda)$ where $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$ denote the matrix
 (2.1) $\left(f_{\alpha}^{[1]}(\Lambda)\right)_{ij}=\begin{cases}\frac{\lambda_{i}^{\alpha}-\lambda_{j}^{\alpha}}{\lambda_{i}-\lambda_{j}}&i\neq j,\\ \alpha\lambda_{i}^{\alpha-1}&i=j.\end{cases}$
-We discuss the origin of this matrix in Section $3.$
+We discuss the origin of this matrix in Section 3.3. Let $\circ$ denote the Hadamard product between two matrices.
 Theorem 2.1 (DLN Geodesic Equations).
-$X(t)=U(t)\Sigma(t)V(t)^{T}$ is a geodesic between $A$ and $B$ on $(\mathfrak{M}_{d},g^{N})$ if it satisfies the Dirichlet boundary value problem
+$X(t)=U(t)\Sigma(t)V(t)^{T}$ is a geodesic between $A$ and $B$ on $(\mathfrak{M}_{d},g^{N})$ if and only if it satisfies the Dirichlet boundary value problem
 (2.2) $\begin{cases}\dot{X}=\mathcal{A}_{N,X}(P),\\ \dot{P}=-U\left(\sum_{p=1}^{N}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]\Sigma+\Sigma\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\right)V^{T},\\ X(0)=A,\\ X(1)=B,\end{cases}$
@@
 Using Riemannian submersion, we are then able to derive explicit solutions to these equations for particular endpoints.
-As an important intermediate result, we characterize the straight lines in $\mathcal{M}$.
-Lemma 2.2 (Characterization of Straight Lines on $\mathcal{M}$).
-Let $\mathbf{A}=\xi(\Lambda,Q_{N},\ldots,Q_{0})$ and $\mathbf{B}=\xi(\tilde{\Lambda},\tilde{Q}_{N},\ldots,\tilde{Q}_{0})$ be the endpoints of interest. Let
+As an important intermediate result, we find a sufficient condition for straight lines in $\mathcal{M}$.
+Let $\mathbf{A}=\xi(\Lambda,Q_{N},\ldots,Q_{0})$ and $\mathbf{B}=\xi(\tilde{\Lambda},\tilde{Q}_{N},\ldots,\tilde{Q}_{0})$ be two endpoints in $\mathcal{M}$. Then, define
 (2.5) $R_{p}=\tilde{Q}_{p}^{T}Q_{p}\in O_{d}.$
-Then, the line
+Lemma 2.2 (Straight Lines on $\mathcal{M}$).
+  1. (1)
+    The line
     (2.6) $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$
-lies in the balanced manifold for all $t\in[0,1]$:
+    is balanced for all $t\in[0,1]$ if and only if
+    (2.7) $(R_{p+1}-R_{p-1})\left(\tilde{\Lambda}^{-1}R_{p}\Lambda\right)^{T},$
+    is skew symmetric for all $1\leq p\leq N-1$.
   2. (2)
-    if $R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$.
-  2. (2)
-    if and only if
-    (2.7) $(R_{p+1}-R_{p-1})M^{T}$
-    are skew symmetric, where
-    (2.8) $M=\tilde{\Lambda}^{-1}R_{p}\Lambda$
-    for all $1\leq p\leq N-1$.
+    If $R_{p}=R_{p-1}$ for all $1\leq p\leq N$, $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in[0,1]$.
 We are now ready to state the second main result. Let $A=U\Sigma V^{T}$ and $B=\tilde{U}\tilde{\Sigma}\tilde{V}^{T}$ be the endpoints in $\mathfrak{M}_{d}$ of the geodesic.
@@
 Remark 2.4.
-Examples of when $A$ and $B$ satisfy this right-left rotation condition is if they are both symmetric or if they have the same singular vectors ($Q=\text{Id}$).
+Examples of when $A$ and $B$ satisfy this right-left rotation condition are if they are both symmetric or if they have the same singular vectors ($Q=\text{Id}$).
 Remark 2.5.
@@
 (2.9) $\begin{split}X(t)&=\left((1-t)A^{1/N}+tB^{1/N}\right)^{N}\\ &=\sum_{p=0}^{N}\binom{N}{p}(1-t)^{p}t^{N-p}A^{p/N}B^{(N-p)/N}.\end{split}$
-This simplification can be shown by applying the theorem and expanding the geodesic formulas using that diagonal matrices commute.
-The limiting behavior as $N\to\infty$ (recognizing $X(t)$ is the expectation of $A^{S}B^{1-S}$ where $S$ is the average of NN Bernoulli(1t)(1-t) random variables) is
+This simplification can be shown by applying the theorem and expanding the geodesic formulas using the commutativity of diagonal matrices.
+The limiting behavior as $N\to\infty$ (recognizing $X(t)$ as the expectation of $A^{S}B^{1-S}$ where $S$ is the average of NN Bernoulli(1t)(1-t) random variables) is
 (2.10) $X(t)=A^{1-t}B^{t}.$
-We discuss this striking simplification more in Appendix A.
-The two important tools we will use to prove Theorem 2.3 in Section 3.4 are (1) the characterization of straight lines on $\mathcal{M}$ (Lemma 2.2) and (2) an orthonormal basis for the horizontal space in $\mathcal{M}$ constructed by [10, 11, Theorem 34] (Lemma 3.10 and see [9] for an exposition).
+The two important tools we will use to prove Theorem 2.3 in Section 3.4 are (1) the sufficient condition for straight lines on $\mathcal{M}$ (Lemma 2.2) and (2) an orthonormal basis for the horizontal space in $\mathcal{M}$ (Lemma 3.11 and [10, 13]).
 Straight lines are automatically geodesics on $(\mathcal{M},\iota)$ because $\iota$ is induced from the Euclidean metric on $\mathbb{M}_{d}^{N}$.
@@
 3.1. Geodesics Review
+For more details on geodesics, we refer the reader to [7]. Recall the following standard definitions and results.
 Definition 3.1 (Geodesic).
-A geodesic between $A,B\in\mathcal{M}$ on $(\mathcal{M},g)$ is an extrema of the action
-(3.1) $\mathcal{A}[x]=\int_{0}^{1}g_{x(t)}(\dot{x}(t),\dot{x}(t))dt$
+A geodesic between $A,B\in\mathcal{X}$ on the Riemannian manifold $(\mathcal{X},g)$ is an extrema of the action
+(3.1) $\mathcal{A}[x]=\int_{0}^{1}L(x,\dot{x})dt,\quad L(x,\dot{x})=g_{x(t)}(\dot{x}(t),\dot{x}(t))$
 over the set of paths
-(3.2) $\mathcal{S}=\{x:[0,1]\to\mathcal{M}|x(0)=A,x(1)=B,x\in\mathcal{C}^{1}([0,1])\}.$
-The minimizing $x$ is known to satisfy the Euler-Lagrange equations
+(3.2) $\mathcal{S}=\{x:[0,1]\to\mathcal{X}|x(0)=A,x(1)=B,x\in\mathcal{C}^{1}([0,1])\}.$
+$L$ is referred to as the Lagrangian.
+Theorem 3.2 (Euler-Lagrange and Hamiltonian Equations).
+A geodesic $x$ satisfies the Euler-Lagrange equations:
 (3.3) $\frac{d}{dt}\frac{\partial L(x,\dot{x})}{\partial\dot{x}}=-\frac{\partial L(x,\dot{x})}{\partial x}.$
-The dual of the Lagrangian optimization problem is the Hamiltonian formulation that reveals an equivalent set of equations of motion.
-Given a Lagrangian $L(x,\dot{x})$, the Hamiltonian is the Legendre transform of $L(x,\cdot)$
-(3.4) $H(x,p)$ $=\sup_{\dot{x}}(\langle p,\dot{x}\rangle-L(x,\dot{x})).$
-It is not hard to see that the change of variables is $p=\frac{\partial L}{\partial\dot{x}}$ when $\dot{x}$ is maximizing. The corresponding canonical equations to the Euler-Lagrange equations are
+and the dual Hamiltonian equations
 (3.4) $\dot{x}$ $=\partial_{p}H(x,p),$
 (3.5) $\dot{p}$ $=-\partial_{x}H(x,p).$
+where $H$ is the Legendre transform of $L(x,\dot{x})$
+(3.6) $H(x,p)=\sup_{\dot{x}}(\langle p,\dot{x}\rangle-L(x,\dot{x})).$
+with change of variables $p=\frac{\partial L}{\partial\dot{x}}$.
+For a Riemannian manifold $(\mathcal{X},g)$, the Hamiltonian $H$ associated with Lagrangian $L(x,\dot{x})=\frac{1}{2}g(\dot{x},\dot{x})$ is always $H(x,p)=\frac{1}{2}g^{-1}(p,p)$, where $g^{-1}$ is the inverse metric tensor acting on cotangent vectors.
 3.2. The Bures-Wasserstein Geometry
-The Bures-Wasserstein geometry [5] is a special case of the DLN model when $N=2$ and $W_{1}=W_{2}^{T}$.
-The parameter space becomes $\mathfrak{M}_{d}$ whereas the observable space is $\mathbb{P}_{d}$.
+The Bures-Wasserstein geometry [4] is a special case of the DLN model when $N=2$ and $W_{1}=W_{2}^{T}$ that will serve as a motivating example for the derivation of Theorem 2.1.
+The parameter space is $\mathfrak{M}_{d}$ whereas the observable space is $\mathbb{P}_{d}$.
 The map $\pi:\mathfrak{M}_{d}\to\mathbb{P}_{d}$ is a Riemannian submersion that maps
@@
 (3.8) $\mathcal{L}^{-1}_{X}(Z)X+X\mathcal{L}^{-1}_{X}(Z)=Z.$
-It is also clear that $\mathcal{L}_{X}(Y)=XY+YX$. An important quantity in the Bures-Wasserstein geometry is the geometric mean of two matrices.
+The inverse metric tensor is $\mathcal{L}_{X}(Y)=XY+YX$. An important quantity in the Bures-Wasserstein geometry is the geometric mean of two matrices.
 Definition 3.3 (Geometric Mean).
@@
 (3.9) $A\#B=A^{1/2}(A^{-1/2}BA^{-1/2})^{1/2}A^{1/2}.$
-Lemma 3.3 (Geometric Mean Properties).
 We will use two key facts about the geometric mean, both of which can be proven directly from the definition.
+Lemma 3.4.
   1. (1)
@@
     $A(A^{-1}\#B)=(AB)^{1/2}$ and $(A^{-1}\#B)A=(BA)^{1/2}$.
-As aforementioned, the Bures-Wasserstein geometry acts as a motivating example for our results.
-Specifically, we present a derivation of the explicit formulas for geodesics by solving the Hamiltonian equations of motion directly that we will generalize to the DLN (although we will not solve the equations explicitly).
+Our main computation in the Bures-Wasserstein geometry is a derivation of the explicit formulas for geodesics by solving the Hamiltonian equations of motion directly.
 Theorem 3.5 (Bures-Wasserstein Geodesics).
@@
 Proof of Theorem 3.5.
-To begin, recall that $T_{X}\mathbb{P}_{d}\cong\text{Symm}_{d}$ and denote $Y=\dot{X}$.
-In order to construct the change of variable from Lagrangian into Hamiltonian coordinates, we must first compute $D_{Y}[L(X,Y)](H)$ where $H\in T_{X}\mathbb{P}_{d}$.
-$D_{Y}\left[\frac{1}{2}\operatorname{Tr}(Y^{T}\mathcal{L}^{-1}_{X}(Y))\right](H)$ $=\frac{1}{2}\operatorname{Tr}(D_{Y}[Y^{T}\mathcal{L}^{-1}_{X}(Y)](H))$
-$=\frac{1}{2}\operatorname{Tr}(H^{T}\mathcal{L}^{-1}_{X}(Y)+Y^{T}\mathcal{L}^{-1}_{X}(H))$
-$=\frac{1}{2}(g_{X}(H,Y)+g_{X}(Y,H))$
-(3.13) $=\operatorname{Tr}(\mathcal{L}^{-1}_{X}(Y)H)$
-where the second to last equality holds because $H$ and $Y$ are in $T_{X}\mathbb{P}_{d}$ and the last equality follows from the cyclic property of trace.
-Thus, we obtain that $P=\mathcal{L}^{-1}_{X}(Y)$ and that $Y=\mathcal{L}_{X}(P)$.
-Now, we can compute the Hamiltonian using the cyclic property of trace and the definition of $P$ and $\mathcal{L}_{X}$ as
-$H(X,P)$ $=\operatorname{Tr}(P^{T}\mathcal{L}_{X}(P))-\frac{1}{2}\operatorname{Tr}(\mathcal{L}_{X}(P)\mathcal{L}^{-1}_{X}(\mathcal{L}_{X}(P)))$
-(3.14) $=\frac{1}{2}\operatorname{Tr}(P\mathcal{L}_{X}(P))=\operatorname{Tr}(P^{2}X).$
+To begin, recall that $T_{X}\mathbb{P}_{d}\cong\text{Symm}_{d}$ and denote $Y=\dot{X}$. We consider the Hamiltonian system, which from the inverse metric tensor $\mathcal{L}$ is
+(3.13) $H(X,P)$ $=\frac{1}{2}\operatorname{Tr}(P\mathcal{L}_{X}(P))=\operatorname{Tr}(P^{2}X).$
 The system of differential equations is now clear from the form of $H(X,P)$. We can solve them to find exact formulas for $X$ and $P$. In particular, $\dot{P}=-P^{2}$ implies
@@
 (3.28) $=(1-t)^{2}A+t^{2}B+t(1-t)\left((AB)^{1/2}+(BA)^{1/2}\right).$
+The uniqueness of this solution also follows from Picard’s theorem, as the right hand sides of Equation 3.11 are smooth. ∎
 3.3. General Equations from Hamiltonian
 We will now generalize the preceding calculation to the DLN.
-First, recall functional calculus on symmetric matrices [4, Chapter 2].
+First, recall functional calculus on symmetric matrices [5, Chapter 2].
 If $S=Q\Lambda Q^{T}$ is a symmetric matrix and $f:\mathbb{R}\to\mathbb{R}$ is a 1D functional, $f$ can be extended to the symmetric matrices as (abusing notation)
@@
 (3.31) $f^{[1]}(\Lambda)_{ij}=\begin{cases}\frac{f(\lambda_{i})-f(\lambda_{j})}{\lambda_{i}-\lambda_{j}}&i\neq j\\ \lim_{\lambda_{i}\to\lambda_{j}}\frac{f(\lambda_{i})-f(\lambda_{j})}{\lambda_{i}-\lambda_{j}}&i=j.\end{cases}$
-$f^{[1]}_{\alpha}$ (Equation 2.1) now clearly comes from Equation 3.32 if $f(x)=x^{\alpha}$. We also have the following derivative computation.
+Equation 2.1 for $f^{[1]}_{\alpha}$ is the special case of Equation 3.31 when $f(x)=x^{\alpha}$. We also have the following derivative computation.
 Lemma 3.6 (Derivatives of $(XX^{T})^{\alpha}$).
@@
 (3.35) $D_{X}\left[XX^{T}\right](Z)=XZ^{T}+ZX^{T}.$
-Rewriting $Z$ as $\tilde{Z}$ in the singular vector coordinates via the transformation $Z=U\tilde{Z}V^{T}$ nets the final simplification that gives the form in the Lemma statement. ∎
+The final simplification comes from expanding using $Z=U\tilde{Z}V^{T}$. ∎
 We also prepare another intermediate result that will be useful for algebraic manipulation.
@@
 Let $K$ and $M$ be symmetric matrices, $\Sigma$ be a diagonal matrix, and $A$ be an arbitrary matrix. Then, we have the following identities.
-(3.37) $\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\operatorname{Tr}([K\circ(A\Sigma)]M)=\operatorname{Tr}(\Sigma[K\circ M]A),$
-(3.38) $\operatorname{Tr}([K\circ(\Sigma A)]M)$ $=\operatorname{Tr}([K\circ(A^{T}\Sigma)]M)=\operatorname{Tr}([K\circ M]\Sigma A).$
+(3.36) $\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\operatorname{Tr}([K\circ(A\Sigma)]M)=\operatorname{Tr}(\Sigma[K\circ M]A)$
 Proof.
-Again, we will just prove for the first equation and note the second follows from almost identical steps.
-Let’s first prove that the first expression is equal to the last.
-We will approach by expanding into coordinates, following Einstein notation convention where repeated indices are summed.
-$\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=K_{ij}\sigma_{i}(A^{T})_{ij}M_{ji}$
-$=\sigma_{i}K_{ij}M_{ji}A_{ji}$
-$=\sigma_{i}K_{ij}M_{ij}A_{ji}$
+First, we prove that the first expression is equal to the last. We will approach by expanding into coordinates.
+$\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\sum_{i}\sum_{j}K_{ij}\sigma_{i}(A^{T})_{ij}M_{ji}$
+$=\sum_{i}\sum_{j}\sigma_{i}K_{ij}M_{ji}A_{ji}$
+$=\sum_{i}\sum_{j}\sigma_{i}K_{ij}M_{ij}A_{ji}$
 (3.37) $=\operatorname{Tr}(\Sigma[K\circ M]A).$
 Similarly, we can prove the equality of the second expression and the third expression using the symmetry of $K$.
-$\operatorname{Tr}([K\circ(A\Sigma)]M)$ $=K_{ij}A_{ij}\sigma_{j}M_{ji}$
-$=\sigma_{j}K_{ij}M_{ji}A_{ij}$
-$=\sigma_{j}K_{ji}M_{ji}A_{ij}$
+$\operatorname{Tr}([K\circ(A\Sigma)]M)$ $=\sum_{i}\sum_{j}K_{ij}A_{ij}\sigma_{j}M_{ji}$
+$=\sum_{i}\sum_{j}\sigma_{j}K_{ij}M_{ji}A_{ij}$
+$=\sum_{i}\sum_{j}\sigma_{j}K_{ji}M_{ji}A_{ij}$
 (3.38) $=\operatorname{Tr}(\Sigma[K\circ M]A).$
-Following similar steps to the Bures Wasserstein case, we can find the Hamiltonian on the manifold $(\mathfrak{M}_{d},g^{N})$ to be
+The Hamiltonian on the manifold $(\mathfrak{M}_{d},g^{N})$ is again given by the inverse metric tensor:
 (3.39) $H(X,P)=\frac{1}{2}\operatorname{Tr}\left(P^{T}\mathcal{A}_{N,X}(P)\right).$
@@
 Proof of Theorem 2.1.
-This first differential equation follows from exactly the same computation as the Bures-Wasserstein case (Equation 3.2) but with $\mathcal{L}^{-1}$ replaced by $\mathcal{A}^{-1}_{N,X}$.
-The only nontrivial equation is the second equation and the associated computation of $D_{X}[H(X,P)](Z)$.
-We begin by applying Lemma 3.5.
-(3.42) $D_{X}[H(X,P)](Z)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}\left(T_{U}+T_{V}\right),$
+The first equation follows directly from the Hamiltonian. The second equation requires computation of $D_{X}[H(X,P)](Z)$. We begin by applying Lemma 3.6.
+(3.40) $D_{X}[H(X,P)](Z)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}\left(T_{U}+T_{V}\right)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}(T_{U})+\operatorname{Tr}(T_{V}),$
 (3.41) $T_{U}=P^{T}U\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]U^{T}PV\Sigma^{\frac{2(p-1)}{N}}V^{T},$
 (3.42) $T_{V}=P^{T}U\Sigma^{\frac{2(N-p)}{N}}U^{T}PV\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right]V^{T}.$
-To clean up further, we also write $P$ in terms of the singular value coordinates by setting $\tilde{P}=U^{T}PV$, leaving just
-(3.45) $T_{U}$ $=\tilde{P}^{T}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]\tilde{P}\Sigma^{\frac{2(p-1)}{N}}$
-(3.46) $T_{V}$ $=\tilde{P}^{T}\Sigma^{\frac{2(N-p)}{N}}\tilde{P}\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right].$
-Applying the trace identities from Lemma 3.6 to Equations 3.45 and 3.46, where we note $f^{[1]}_{\alpha}(\Sigma^{2})$ is symmetric, and substituting in $M_{1}$ and $M_{2}$ from Equation 2.3 reveal
+To clean up further, we can write $P$ in terms of the singular value coordinates by setting $\tilde{P}=U^{T}PV$ and using the cyclic property of trace to see that
+(3.43) $\operatorname{Tr}(T_{U})$ $=\operatorname{Tr}\left(\tilde{P}^{T}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]\tilde{P}\Sigma^{\frac{2(p-1)}{N}}\right)$
+(3.44) $\operatorname{Tr}(T_{V})$ $=\operatorname{Tr}\left(\tilde{P}^{T}\Sigma^{\frac{2(N-p)}{N}}\tilde{P}\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right]\right).$
+Applying the identities from Lemma 3.7 to Equations 3.43 and 3.44 ($f^{[1]}_{\alpha}(\Sigma^{2})$ is symmetric) and substituting in $M_{1}$ and $M_{2}$ from Equation 2.3 shows
 […]
 […]
-(3.49) $=\operatorname{Tr}\left(V\sum_{p=1}^{N}\left(\Sigma\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]+\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\Sigma\right)U^{T}Z\right)$
+(3.47) $=\operatorname{Tr}\left(V\sum_{p=1}^{N}\left(\Sigma\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]+\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\Sigma\right)U^{T}Z\right).$
 This lets us conclude that the second canonical equation in the Hamiltonian formulation is
@@
 Definition 3.8 (Vertical and Horizontal Spaces).
-Let $(\mathcal{M},g)$ and ($\mathcal{N},h)$ be linked by a differentiable map $\pi:\mathcal{M}\to\mathcal{N}$.
-Then, there exists a direct sum decomposition of the tangent space at any point $m\in\mathcal{M}$ via the differential of $\phi$.
-(3.51) $T_{m}\mathcal{M}=\ker D\pi[m]\oplus(\ker D\pi[m])^{\perp}=\mathcal{V}_{m}\oplus\mathcal{H}_{m},$
+Let $(\mathcal{X},g)$ and ($\mathcal{Y},h)$ be linked by a differentiable map $\pi:\mathcal{X}\to\mathcal{Y}$.
+Then, there exists a direct sum decomposition of the tangent space at any point $m\in\mathcal{X}$ via the differential of $\phi$.
+(3.49) $T_{m}\mathcal{X}=\ker D\pi[m]\oplus(\ker D\pi[m])^{\perp}=\mathcal{V}_{m}\oplus\mathcal{H}_{m},$
 where $\mathcal{V}_{m}$ and $\mathcal{H}_{m}$ denote the vertical and horizontal spaces at $m$ respectively.
 Definition 3.9 (Riemannian Submersion).
-Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be two Riemannian manifolds and let $T_{m}\mathcal{M}=\mathcal{V}_{m}\oplus\mathcal{H}_{m}$ be the direct sum decomposition of the tangent space at $m\in\mathcal{M}$.
-A mapping $\pi:\mathcal{M}\to\mathcal{N}$ is a Riemannian submersion if and only if the differential $D\pi[m]:T_{m}\mathcal{M}\to T_{\pi(m)}\mathcal{N}$ satisfies the following for all $m\in\mathcal{M}$:
+Let $(\mathcal{X},g)$ and $(\mathcal{Y},h)$ be two Riemannian manifolds and let $T_{m}\mathcal{X}=\mathcal{V}_{m}\oplus\mathcal{H}_{m}$ be the direct sum decomposition of the tangent space at $m\in\mathcal{X}$.
+A mapping $\pi:\mathcal{X}\to\mathcal{Y}$ is a Riemannian submersion if and only if the differential $D\pi[m]:T_{m}\mathcal{X}\to T_{\pi(m)}\mathcal{Y}$ satisfies the following for all $m\in\mathcal{X}$:
   1. (1)
@@
     $D\pi[m]$ restricted to the horizontal space is an isometry i.e. $\|z\|_{g}=\|D\pi[m](z)\|_{h}$ for all $z\in\mathcal{H}_{m}$.
-A core fact underlying [5, Section 4] and Theorem 2.3 is that understanding the horizontal space of a submersion allows us to identify geodesics that are preserved under the submersion.
+A core fact underlying [4, Section 4] and Theorem 2.3 is that understanding the horizontal space of a submersion allows us to identify geodesics that are preserved under the submersion, as described by the following theorem [12, 4].
 Theorem 3.10 (Submersions Preserve Horizontal Geodesics).
-Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be Riemannian manifolds and $\phi:\mathcal{M}\to\mathcal{N}$ be a Riemannian submersion.
-Let $x:[0,1]\to\mathcal{M}$ be a geodesic between $A,B\in\mathcal{M}$ in the geometry $(\mathcal{M},g)$ so that $x^{\prime}(0)\in\mathcal{H}_{A}$ i.e.
+Let $(\mathcal{X},g)$ and $(\mathcal{Y},h)$ be Riemannian manifolds and $\phi:\mathcal{X}\to\mathcal{Y}$ be a Riemannian submersion.
+Let $x:[0,1]\to\mathcal{X}$ be a geodesic between $A,B\in\mathcal{X}$ in the geometry $(\mathcal{X},g)$ so that $x^{\prime}(0)\in\mathcal{H}_{A}$ i.e.
 is horizontal.
@@
   2. (2)
-    $\phi\circ x$ is a geodesic in $(\mathcal{N},h)$ between $\phi(A)$ and $\phi(B)$ and is the same length as $x$.
-Thus, if we are able to classify horizontal geodesics in $(\mathcal{M},\iota)$, Theorem 3.9 will automatically give explicit formulas for geodesics in $(\mathfrak{M}_{d},g^{N})$ through $\phi$, the submersion.
+    $\phi\circ x$ is a geodesic in $(\mathcal{Y},h)$ between $\phi(A)$ and $\phi(B)$ and is the same length as $x$.
+Thus, if we are able to classify horizontal geodesics in $(\mathcal{M},\iota)$, Theorem 3.10 will give explicit formulas for geodesics in $(\mathfrak{M}_{d},g^{N})$ through the submersion $\phi$.
 The simplest class of geodesics in $(\mathcal{M},\iota)$ are straight lines.
-Thus, we begin with a proof of Lemma 2.2 that characterizes the straight lines in $\mathcal{M}$.
+Thus, we begin with a proof of Lemma 2.2 that describes a set of straight lines in $\mathcal{M}$.
 Proof of Lemma 2.2.
-We wish to find the condition that $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$ is in $\mathcal{M}$ for all $t\in(0,1)$ if $\mathbf{A},\mathbf{B}\in\mathcal{M}$.
+First, we wish to find the equivalent condition to $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$ being balanced for all $t\in(0,1)$ when $\mathbf{A},\mathbf{B}\in\mathcal{M}$.
 Fix $p\in\{1,2,\ldots,N-1\}$.
-We would like to show that
+We would like to find an equivalent condition to
 (3.50) $W_{p+1}(t)^{T}W_{p+1}(t)=W_{p}(t)W_{p}(t)^{T},$
 where $W_{p}(t)=(1-t)W_{p}(0)+tW_{p}(1)$ for all $p$.
-Expanding the algebra out and using that $W_{p}(0)=A_{p}$ and $W_{p}(1)=B_{p}$ are both in the balanced manifold, we find that $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in(0,1)$ if and only if
+Expanding the algebra out and using that $W_{p}(0)=A_{p}$ and $W_{p}(1)=B_{p}$ are both in the balanced manifold, we find that $\mathbf{W}(t)$ is balanced for all $t\in(0,1)$ if and only if
 (3.51) $B_{p+1}^{T}A_{p+1}+A_{p+1}^{T}B_{p+1}=B_{p}A_{p}^{T}+A_{p}B_{p}^{T}.$
@@
 (3.53) $\tilde{\Lambda}(R_{p+1}-R_{p-1})\Lambda$ $=-R_{p}\Lambda(R_{p+1}-R_{p-1})^{T}\tilde{\Lambda}R_{p}.$
-The sufficient condition ($R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$) clearly results in this condition being satisfied for all $t\in(0,1)$.
-From here, it is a simple rewriting in terms of $M$ to get the final equivalent condition:
-(3.56) $R_{p+1}-R_{p-1}$ $=-M(Q_{p+1}-Q_{p-1})^{T}M^{-T}.$
-(3.57) $(R_{p+1}-R_{p-1})M^{T}$ $=-M(Q_{p+1}-Q_{p-1})^{T}.$
-The structure demonstrates that the difference
-(3.58) $R_{p+1}-R_{p-1}$
-behaves like a skew symmetric matrix after changing into coordinates given by $M$ (or essentially $R_{p}$). ∎
-We now state an orthonormal basis for the horizontal space in $\mathcal{M}$ computed in [10, Section 6] and [11, Theorem 34]. Let
+We can rewrite in terms of $M=\tilde{\Lambda}^{-1}R_{p}\Lambda$ to get the final equivalent condition:
+(3.54) $R_{p+1}-R_{p-1}$ $=-M(R_{p+1}-R_{p-1})^{T}M^{-T}.$
+(3.55) $(R_{p+1}-R_{p-1})M^{T}$ $=-M(R_{p+1}-R_{p-1})^{T}.$
+The structure demonstrates that $\mathbf{W}(t)$ is balanced if and only if the difference
+(3.56) $(R_{p+1}-R_{p-1})M^{T}$
+is a skew symmetric matrix.
+It remains to show that if $R_{p}=R_{p-1}$ for all $1\leq p\leq N$, $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in[0,1]$.
+In this case, the above skew symmetry condition is immediate as all differences $R_{p+1}-R_{p-1}$ are 0, implying that $\mathbf{W}(t)$ is balanced for all $t\in[0,1]$.
+Thus, we just need to show that the coordinates of $\mathbf{W}(t)$ stay full rank at all intermediate $t$.
+Fix $1\leq p\leq N$ and consider the $p$-th coordinate of $\mathbf{W}(t)$. Since $R_{p}$ is the same for all $p$, we drop the index for clarity.
+(3.57) $\mathbf{W}_{p}(t)$ $=(1-t)A_{p}+tB_{p}$
+(3.58) $=(1-t)Q_{p}\Lambda Q_{p-1}^{T}+t\tilde{Q}_{p}\tilde{\Lambda}\tilde{Q}_{p-1}^{T}$
+(3.59) $=(1-t)\tilde{Q}_{p}R\Lambda R^{T}\tilde{Q}_{p-1}^{T}+t\tilde{Q}_{p}\tilde{\Lambda}\tilde{Q}_{p-1}^{T}$
+(3.60) $=\tilde{Q}_{p}\left((1-t)R\Lambda R^{T}+t\tilde{\Lambda}\right)\tilde{Q}_{p-1}^{T}.$
+Because $\mathbb{P}_{d}$ is convex, the interior convex combination is also in $\mathbb{P}_{d}$ as $R\Lambda R^{T}\in\mathbb{P}_{d}$.
+Since $\tilde{Q}_{p},\tilde{Q}_{p-1}\in O_{d}$ and are invertible, we conclude that $\mathbf{W}_{p}(t)$ is full rank for all $t\in[0,1]$ and $\mathbf{W}(t)\in\mathcal{M}$ follows.
+∎
+We now state a basis for the horizontal space in $\mathcal{M}$ computed in [10, Section 6] and [13, Theorem 34]. Let
 (3.61) $\mathbf{M}\in\mathcal{M},\mathbf{M}=\left(Q_{N}\Lambda^{\frac{1}{N}}Q_{N-1}^{T},Q_{N-1}\Lambda^{\frac{1}{N}}Q_{N-2}^{T},\dots,Q_{1}\Lambda^{\frac{1}{N}}Q_{0}^{T}\right)$
 so that $\mathbf{M}=\xi(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{1})$ and $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$.
-Furthermore, let $q_{p,i}$ denote the $i$th column of $Q_{p}$.
-Then, define the following special rank one elements of $\mathbb{M}_{d}^{N}$:
-(3.60) $\begin{cases}\mathbf{l}^{k}=\left(l^{k}_{N},l^{k}_{N-1},\ldots,l^{k}_{1}\right),\\ l^{k}_{s}=\frac{1}{\sqrt{N}}q_{s,k}q_{s-1,k}^{T},\end{cases}$
-(3.61) $\begin{cases}\mathbf{u}^{k,l,0}=\left(u_{N}^{k,l,0},u_{N-1}^{k,l,0},\ldots,u_{1}^{k,l,0}\right),\\ u^{k,l,0}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{s-1}\lambda_{l}^{N-s}q_{s,l}q_{s-1,k}^{T},\end{cases}$
-and
-(3.62) $\begin{cases}\mathbf{u}^{k,l,N}=\left(u_{N}^{k,l,N},u_{N-1}^{k,l,N},\ldots,u_{1}^{k,l,N}\right),\\ u^{k,l,N}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{N-s}\lambda_{l}^{s-1}q_{s,k}q_{s-1,l}^{T}.\end{cases}$
-Lemma 3.10 (Orthonormal Basis for Horizontal Space in $\mathcal{M}$).
-(3.63) $\begin{cases}\mathbf{l}=\{l_{k}|1\leq k\leq d\},\\ \mathbf{u}^{0}=\{u^{k,l,0}|1\leq k<l\leq d\},\\ \mathbf{u}^{N}=\{u^{k,l,N}|1\leq k<l\leq d\}\end{cases}$
-form a basis for $\mathcal{H}_{\mathbf{M}}=(\ker\phi^{*}[\mathbf{M}])^{\perp}$.
-Remark 3.11.
-Note a matrix $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{M}}$ (span of $\mathbf{l}$, $\mathbf{u}^{k,l,0}$, and $\mathbf{u}^{k,l,N}$) if and only if there exists a set of coefficients $\alpha\in\mathbb{M}_{d}$ such that
+The horizontal space at $\mathbf{M}$ is $\mathcal{H}_{\mathbf{M}}=\ker(\phi^{*}[\mathbf{M}])^{\perp}$.
+Let $q_{p,i}$ denote the $i$-th column of $Q_{p}$.
+Lemma 3.11 (Basis for Horizontal Space in $\mathcal{M}$).
+Let $Z$ be an arbitrary matrix in $\mathbb{M}_{d}^{N}$. Then, $Z\in\mathcal{H}_{\mathbf{M}}$ if and only if there exists a set of coefficients $\alpha\in\mathbb{M}_{d}$ such that
 (3.62) $Q_{p}\alpha Q_{p-1}^{T}=Z_{p}$
-for all $1\leq p\leq N$.
-Now that we understand both the straight lines and horizontal space in $\mathcal{M}$, we can use $\phi$ and Theorem 3.9 to construct new explicit formulas for geodesics on $\mathfrak{M}_{d}$.
+for all $1\leq p\leq N$. Equivalently, $\mathcal{H}_{\mathbf{M}}$ is spanned by
+(3.63) $\left\{\left(q_{N,k}q_{N-1,l}^{T},q_{N-1,k}q_{N-2,l}^{T},\ldots q_{1,k}q_{0,l}^{T}\right)|1\leq k,l\leq d\right\}.$
+We can now use $\phi$ and Theorem 3.10 to construct new explicit formulas for geodesics on $\mathfrak{M}_{d}$.
 Proof of Theorem 2.3.
@@
 Clearly, $\phi(\mathbf{A})=A$.
-Furthermore, Lemma 3.10 gives a set of orthonormal basis vectors for the horizontal space $\mathcal{H}_{\mathbf{A}}$:
-(3.68) $l^{k}_{s}$ $=\frac{1}{\sqrt{N}}\begin{cases}u_{k}e_{k}^{T}&s=N\\ e_{k}v_{k}^{T}&s=1\\ e_{k}e_{k}^{T}&\text{o/w}\end{cases},$
-(3.69) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{s-1}{N}}\sigma_{l}^{\frac{N-s}{N}}\begin{cases}u_{l}e_{k}^{T}&s=N\\ e_{l}v_{k}^{T}&s=1\\ e_{l}e_{k}^{T}&\text{o/w}\end{cases},$
-(3.70) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{N-s}{N}}\sigma_{l}^{\frac{s-1}{N}}\begin{cases}u_{k}e_{l}^{T}&s=N\\ e_{k}v_{l}^{T}&s=1\\ e_{k}e_{l}^{T}&\text{o/w}\end{cases}.$
-By Remark 3.11, $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{A}}$ if and only if there exists $\alpha\in\mathbb{M}_{d}$ such that
+Furthermore, by Lemma 3.11, $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{A}}$ if and only if there exists $\alpha\in\mathbb{M}_{d}$ such that
 (3.67) $U\alpha$ $=Z_{N},$
@@
 (3.69) $\alpha V^{T}$ $=Z_{1}.$
-It is immediately obvious that $\mathbf{A}$ is in its own horizontal space $\mathcal{H}_{\mathbf{A}}$ using the coefficient matrix $\alpha=\Sigma^{\frac{1}{N}}$.
+It is immediate that $\mathbf{A}$ is in its own horizontal space $\mathcal{H}_{\mathbf{A}}$ using the coefficient matrix $\alpha=\Sigma^{\frac{1}{N}}$.
 Now, also consider
@@
 Clearly, $\phi(\mathbf{B})=B$.
-Furthermore, $\mathbf{A}$ and $\mathbf{B}$ satisfy the sufficient condition of Lemma 2.2 since $Q^{T}=\tilde{U}^{T}U=\tilde{V}^{T}V$ and $R_{p}$ is constant for all $p$.
-Thus, the line $(1-t)\mathbf{A}+t\mathbf{B}$ is indeed the geodesic between $\mathbf{A}$ and $\mathbf{B}$ in $(\mathcal{M},\iota)$.
+Furthermore, $\mathbf{A}$ and $\mathbf{B}$ satisfy the sufficient condition of Lemma 2.2 since $Q^{T}=\tilde{U}^{T}U=\tilde{V}^{T}V$ and $R_{p}$ is constant for all $p$, so we conclude that $x(t)\in\mathcal{M}$ for all $t\in[0,1]$.
+Thus, the line $(1-t)\mathbf{A}+t\mathbf{B}$ is indeed the geodesic between $\mathbf{A}$ and $\mathbf{B}$ in $(\mathcal{M},\iota)$ because $\iota$ is flat.
 Now we will show $x^{\prime}(0)$ is horizontal for the chosen $\mathbf{A}$ and $\mathbf{B}$.
 We have already shown that $\mathbf{A}$ is in $\mathcal{H}_{\mathbf{A}}$.
-By Equation 3.66, it is now sufficient to show that $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ as well.
-We thus return to solving the system of equations in Remark 3.11 with $\mathbf{Z}=\mathbf{B}$:
+By Equation 3.65, it is enough to show that $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ as well.
+Again, Lemma 3.11 provides a sufficient system of equations to solve with $\mathbf{Z}=\mathbf{B}$:
 (3.71) $U\alpha$ $=\tilde{U}\tilde{\Sigma}^{\frac{1}{N}}Q^{T},$
 (3.72) $\alpha$ $=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T},$
-(3.77) $\alpha V^{T}$ $=Q\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T},$
-and has the obvious solution given by the middle equations of $\alpha=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T}$ which is consistent with the others by definition of $Q$.
+(3.73) $\alpha V^{T}$ $=Q\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T}.$
+The solution given by the middle equations of $\alpha=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T}$ can be checked as consistent with the others by definition of $Q$.
 Thus, $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$.
 Since we have shown that both $\mathbf{B}$ and $\mathbf{A}$ are in $\mathcal{H}_{\mathbf{A}}$ and that the geodesic linking them is a straight line, it follows from Equation 3.65 that $x^{\prime}(0)$ is horizontal.
-Applying Theorem 3.9, we conclude that $\phi\circ x$ is a geodesic in $(\mathfrak{M}_{d},g^{N})$ of the same length as $x$.
-The explicit formula comes directly from applying $\phi$ to $x$ between our parameterized $\mathbf{A}$ and $\mathbf{B}$.
+Furthermore, note that because all coordinates in $x(t)\in\mathcal{M}$, $\phi\circ x\in\mathfrak{M}_{d}$ for all $t\in[0,1]$ as $\phi\circ x$ becomes a product of invertible matrices.
+Applying Theorem 3.10, we conclude that $\phi\circ x$ is a geodesic in $(\mathfrak{M}_{d},g^{N})$ of the same length as $x$ and the explicit formula comes directly from applying $\phi$ to $x$ between our parameterized $\mathbf{A}$ and $\mathbf{B}$.
 ∎
@@
 The main follow up work is (1) solving the general case geodesic equations for arbitrary endpoints and (2) generalizing geodesic results to low rank (rank <d<d) matrices, not just $\mathcal{M}$, as the Riemannian metric and submersion structure exists in all ranks [3, 10, 13].
-Once geodesics and distances are understood in the geometries underlying the DLN, one can consider reconnecting them to the vector field given by the gradient flow to make conclusions about training dynamics, similar to [8, Section 3.3].
+Once geodesics and distances are understood in the geometries underlying the DLN, one can consider reconnecting them to the vector field given by the gradient flow to make conclusions about training dynamics, similar to [9, Section 3.3].
 Acknowledgements
@@
 This work was partially supported by NSF grant 2407055.
-Appendix A Informal Note on Equation 2.11
-The geodesic in Equation 2.11 also appears in another geometry [4] when the endpoints are positive definite diagonal matrices.
-In particular, we endow $\mathbb{P}_{d}$ with the trace metric: at a matrix $P\in\mathbb{P}_{d}$ and tangent vector $Z\in T_{P}\mathbb{P}_{d}$, we let
-(A.1) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(P^{-1}ZP^{-1}Z).$
-Consider $P$ positive definite and diagonal, and let $Z$ commute with $P$ (e.g. also diagonal). By the properties of trace,
-(A.2) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(ZP^{-2}Z).$
-Between any matrices $P_{0},P_{1}\in\mathbb{P}_{d}$, the geodesic is exactly $P_{0}^{1-t}P_{1}^{t}$.
-Interestingly, $g^{N}$ as $N\to\infty$ coincides with the trace metric under these assumptions.
-Consider a variant of $\mathcal{A}_{N,\cdot}$ with an additional $\frac{1}{N}$ factor in front so that a $N\to\infty$ limit exists.
-Then, indeed, if $P$ is diagonal and commutes with $Z$,
-(A.3) $\mathcal{A}_{N,P}(Z)=\frac{1}{N}\sum_{p=1}^{N}(P^{2})^{\frac{N-p}{N}}Z(P^{2})^{\frac{p-1}{N}}=(P^{2})^{\frac{N-1}{N}}Z.$
-The limit as $N\to\infty$ is just $P^{2}Z$. Thus, $\mathcal{A}^{-1}_{\infty,P}=P^{-2}Z$ and we conclude that
-(A.4) $\|Z\|_{g^{\infty}}=\operatorname{Tr}(Z\mathcal{A}^{-1}_{\infty,P}Z)=\operatorname{Tr}(ZP^{-2}Z)=\|Z\|_{\operatorname{Tr}}.$
-More generally, if $X\in\mathfrak{M}_{d}$ is symmetric, for all $Z\in T_{X}\mathfrak{M}_{d}$ where $X$ and $Z$ commute, we have that this agreement between $g^{\infty}$ and the trace metric.
-Since along the geodesic $X(t)$ between $A$ and $B$ $\dot{X}\in T_{X}\mathfrak{M}_{d}$ is always diagonal, we have agreement between the trace metric and DLN metric for the entire geodesic.
-Contrast this with the smallest nontrivial depth, $N=2$, where if $X$ is symmetric we know already that the DLN metric coincides with the Bures Wasserstein metric for symmetric tangent vectors.